OpenStudy (anonymous):

Simplify (7a^2 – 11a – 9) + (13a^2 + 21a – 3) Answer 20a^2 + 10a – 6 20a^2 – 32a – 6 20a^2 + 10a – 12 20a^2 – 32a – 12 Simplify (7a^2 – 11a – 9) + (13a^2 + 21a – 3) Answer 20a^2 + 10a – 6 20a^2 – 32a – 6 20a^2 + 10a – 12 20a^2 – 32a – 12 @Mathematics

hero (hero):

c

OpenStudy (anonymous):

Add like terms... a^2s with a^2s a's with a's numbers with numbers

OpenStudy (anonymous):

C

OpenStudy (anonymous):

7a^2 + 13a^2 = 20a^2 -11a + 21a = 10a -9 - 3 = -12